OpenStudy (anonymous):

(x^2-25/x^2-6x+5)(x^2-1/x^2-2x+1)= x+5/x-2 please explain cuz I got something a lil different

OpenStudy (anonymous):

Each of these factor, so we have: \[(x-5)(x+5)/(x-1)(x-5) * (x-1)(x+1)/(x-1)(x-1)\] the (x-5)'s cancel out on the first one, and the (x-1)'s cancel out on the 2nd one

OpenStudy (campbell_st):

\[((x - 5)(x +5))/((x - 5)(x+1)) = (x + 5)/(x + 1)\] \[((x + 1)(x-1))/(x - 1)(x -1)) = (x+1)/(x-1)\] (x + 10 will cancel leaving (x+5)/(x-1)

OpenStudy (anonymous):

\[(x^2-25)(x^2-1)=(x+5)(x-5)(x-1)(x+1)\] \[(x^2-6x+5)(x^2-2x+1)=(x-5)(x-1)(x-1)(x-1)\]

OpenStudy (anonymous):

so we have \[(x+5)/(x-1) * (x+1)/(x-1)\]

OpenStudy (anonymous):

@campbell you messed up on your factoring

OpenStudy (anonymous):

Got that does it some how reduce further to get x+5/x-2?

OpenStudy (anonymous):

I wouldnt think so but I could be wrong

OpenStudy (anonymous):

That's what I got too but then the book says the x+5/x-2 answer

OpenStudy (anonymous):

I'm thinking the book got it wrong.

OpenStudy (campbell_st):

x^2 - 2x + 1 is a perfect square (x-1)^2 or (x -1)(x-1)

OpenStudy (anonymous):

Think soo too cuz I know I did the process right and if you came up with the same answer then ya

OpenStudy (anonymous):

@campbell I meant on the first denominator you messed up

OpenStudy (anonymous):

book can be wrong again but wanna make sure.... x^2-x+11 (prime right)

OpenStudy (anonymous):

Yes, thats prime